# Coverage-Guaranteed Speech Emotion Recognition via Calibrated Uncertainty-Adaptive Prediction Sets

### v2

@@ Abstract
Road rage, stemming from emotional suppression and outbursts, can lead to traffic incidents such as rear-end collisions, crashes, and aggressive confrontations, posing significant risks to road safety and public security. Speech Emotion Recognition (SER) methods can detect early signs of negative emotions and provide real-time alerts to drivers, effectively reducing the likelihood of accidents. However, previous studies, such as Hidden Markov Models (HMMs) and Long Short-Term Memory (LSTM) networks, predominantly rely on one-dimensional speech signals, and are prone to overfitting and miscalibration issues, undermining their reliability in high-stakes applications. In this paper, we introduce a risk management framework that produces prediction sets with statistically rigorous guarantees of correctness coverage marginally over the test set. Specifically, we separate a held-out calibration set and develop a binary loss function for each data point to determine whether the ground-truth label is covered by the prediction set, which is calibrated by a data-driven threshold $\lambda$. We develop a joint loss function on the calibration set, and given a user-specified risk level $\alpha$, we then calibrate the threshold $\lambda$ to strictly bound it such that we can manage the expectation of the loss of each test point by this predefined upper bound $\alpha$. We conduct extensive evaluations on 6 base models on 2 datasets. Empirical results demonstrate that we can always bound the average correctness coverage rate (i.e., $\geq 1-\alpha$) at various user-specified risk levels. Moreover, we consistently control the marginal error rate across various splitting ratios (e.g., 0.1) between the calibration and test set, highlighting the robustness and effectiveness of our framework. Furthermore, we propose a small batch online calibration framework based on the local exchangeability assumption, which is suitable for complex scenarios where there is data domain offset or non-IID between batches. By constructing a non-negative test martingale, we ensure the coverage of the prediction set in dynamic scenarios, and verify the effectiveness of this method in controlling the coverage of the prediction set through cross-dataset experiments.
@@ 1 Introduction
In recent years, road rage has become an increasingly serious traffic safety problem worldwide. With the acceleration of urbanization and the increase in car ownership, traffic congestion and driving pressure are increasing, resulting in more and more drivers showing extreme emotional reactions while driving. Studies have shown that road rage not only increases the incidence of traffic accidents, but also exacerbates conflicts between drivers and other road users, further threatening public safety. According to the National Highway Traffic Safety Administration, more than one-third of traffic accidents are related to drivers losing control of their emotions, and this proportion is on the rise.
Speech emotion recognition technology, as a means of emotion recognition based on speech signal analysis, has gradually become an effective tool for solving problems such as road rage. By capturing the driver’s voice characteristics in real time and analyzing his or her emotional state, speech emotion recognition can provide timely emotional feedback to the traffic management system, helping to predict and intervene in potential dangerous driving behaviors. This technology can not only provide more accurate emotion monitoring in smart cars, but can also be applied to other emotion-related fields, such as customer service and mental health monitoring. Therefore, the importance of speech emotion recognition is increasing day by day, becoming one of the core research directions in the fields of intelligent transportation systems, unmanned driving technology, and emotion analysis.
With the rapid development of deep learning technology, speech emotion recognition tasks based on convolutional neural networks have made significant progress. This paper extracts the Mel-spectrogram features of speech signals to characterize different emotion categories and conducts experimental comparisons. Despite this, the model may still encounter overfitting problems during training, resulting in performance degradation on new and unseen data. In addition, current deep learning models generally lack confidence or uncertainty measures for prediction results, which means that in practical applications, the model’s predictions may be unreliable. Recent studies have pointed out that the miscalibration problem still exists. For example, Zhao et al. (2023) pointed out that although convolutional neural networks show high accuracy in speech emotion recognition tasks, the confidence scores of their prediction results are often too high and fail to truly reflect the actual performance of the model, resulting in incorrect emotion recognition. Therefore, improving the uncertainty measure of the model has become the key to improving the performance of speech emotion recognition systems Psaros et al. (2023); Wang et al. (2025a).
To solve these problems, the conformal prediction method provides a new solution. CP can not only give a predicted label, but also build a set of all possible prediction results and provide a confidence guarantee for each prediction result. To ensure that the confidence of the selected label meets the risk tolerance specified by the user, specifically, the Mel-spectrogram feature is used as the feature extraction method in the speech emotion recognition task, and 5 neural networks with basically the same parameter quantity are trained on the IEMOCAP dataset to obtain a pre-trained model. After the feature extraction and model training are completed, the Conformal Prediction framework is used to evaluate the model to ensure the reliability of the prediction results. Specifically, by defining non-conformal scores and associating these scores with the real emotion labels, the calibration set is used to calibrate the model. In this process, the test data is processed according to the risk level (alpha value) specified by the user, and finally a prediction set with marginal coverage guarantee is generated. This process ensures that the emotion recognition model not only gives the emotion label, but also provides a confidence estimate for the label, thereby ensuring the statistical reliability of the prediction results.
Although CP can effectively provide coverage guarantee, it only solves the coverage problem of the prediction set without considering the reliability of the model under different risk levels. In order to solve this problem, a risk control framework is further introduced. The core idea of this framework is to adjust the risk level of the model on the basis of maintaining coverage guarantee to cope with various uncertainties that may arise in practical applications. Under the risk control framework, the beta value is used to quickly determine the optimal prediction set size through binary search. By processing the calibration set, the size of the prediction set can be adjusted under the constraint of a given risk level to meet the needs of risk control. At the same time, the loss function has been improved so that it not only focuses on the prediction error rate, but also considers the size of each prediction set and its uncertainty. The improvement of the loss function enables the framework to be flexibly applied to different tasks and scenarios, such as emotion recognition, intelligent traffic monitoring, etc. By combining CP with risk control, not only the coverage of the model is guaranteed, but also the robustness and accuracy of the model are improved, so that it can adapt to different practical application needs. Experimental results show that the framework can achieve coverage guarantee at different risk levels, while reducing the uncertainty of the model to a certain extent, making it more valuable for practical applications. However, traditional CP methods rely on a strict assumption of data exchangeability, which is valid in controlled environments but often fails in dynamic real-world scenarios, such as speech emotion recognition, where the data distribution may change between batches or domains. Such a non-exchangeable environment leads to a degradation of coverage guarantees, thus limiting the practical application of CP. To overcome this limitation, we propose a novel framework to ensure the coverage of the prediction set in dynamic scenarios by performing conformal prediction via a mini-batch online calibration mechanism. Specifically, we relax the global exchangeability assumption by adopting a local exchangeability assumption within a mini-batch. By constructing a non-negative test martingale, our method ensures coverage even when the distribution changes between batches, such as speaker differences or acoustic environment differences.
In order to verify the effectiveness of the framework, a large number of experiments were carried out. First, starting from the pre-experimental part, the performance of different feature selection methods and pre-trained models was tested. In the pre-experimental stage, a variety of feature extraction methods were compared, including mel-spectrogram features, chromaticity features, and spectral contrast features. Through testing on the TESS dataset, it was found that the performance of mel-spectrogram features was better than other feature extraction methods, achieving a test set accuracy of 99.82%. Therefore, mel-spectrogram features were finally selected as the speech feature extraction method. Next, several different neural network models, including ResNet, MobileNetV3, ShuffleNetV2, etc., were trained on the IEMOCAP dataset, and their performance on emotion recognition tasks was tested. Although the accuracy of these models on the training set reached 100%, the accuracy on the test set dropped to 41.60%, 31.94%, and 34.13%, respectively. This phenomenon indicates that the model has a certain overfitting problem when facing unseen data, so the Conformal Prediction framework is introduced to ensure the reliability and statistical significance of the model prediction results. In order to prove the coverage guarantee of the framework, two rounds of verification were carried out. The first round of experiments was conducted on the same dataset to ensure that the model can provide stable performance on both the calibration set and the test set, verifying that Conformal Prediction can guarantee the accuracy of the prediction results under a given risk level. The second round of experiments conducted cross-dataset validation, that is, the model trained on the IEMOCAP dataset was applied to the TESS dataset to test the generalization ability of the model. The experimental results show that the model can still maintain a high coverage on different datasets and can effectively adapt to different data distributions, verifying the robustness of the framework. Next, the uncertainty of different models is evaluated based on the prediction set size. During the training and testing process of each model, the prediction set size is calculated, and the uncertainty of the model is quantified according to the prediction set size under each alpha value. By conducting multiple experiments under approximately the same parameter conditions, we found that when the uncertainty of the model increases, the prediction set size usually increases. This shows that a larger prediction set means that the model has lower confidence in the recognition of certain emotion categories. To control this uncertainty, we adopt the risk control framework to conduct the same experiment, and combine it with the improved loss function to optimize the prediction set size to ensure that the model can provide reliable prediction results under different risk levels. The final experimental results show that both frameworks can maintain the coverage guarantee and robustness of the final results under different datasets and risk levels, ensuring that the emotion recognition system can meet the needs of practical applications. Crucially, to verify our small batch martingale value method in the scenario of non-IID and domain shift, we designed an additional experiment in which the calibration and test data come from different datasets (IEMOCAP for calibration and TESS for testing), explicitly violating global exchangeability. The data is split into small batches (5 calibration samples + 1 test sample per batch), and the small amount of data within the batch is assumed to be locally exchangeable, while allowing for distribution shifts across batches. The coverage of traditional CP methods that rely on global exchangeability will drop significantly. In contrast, our martingale-based method brings the coverage close to the theoretical bound. This experiment finally proves the effectiveness of our framework in the scenario of non-IID and domain shift.
@@ 2.1 Speech Recognition
The history of speech recognition technology can be traced back to the 1950s. The initial systems relied on traditional algorithms based on template matching, which usually required users to pronounce in a specific way. With the improvement of computing power, hidden Markov models (HMMs) have become the mainstream technology for speech recognition, which can effectively model speech signals in time series. However, these traditional methods still face problems with recognition accuracy and noise interference, especially in complex speech environments. The performance of traditional speech recognition systems often cannot meet the requirements of high accuracy and high robustness. After entering the era of deep learning, speech recognition technology based on neural networks has completely changed this situation. In particular, the introduction of convolutional neural networks (CNNs) and recurrent neural networks (RNNs) has made the processing of speech signals more efficient and accurate. Technologies such as deep neural networks (DNNs) and long short-term memory networks (LSTMs) have greatly improved the accuracy and real-time performance of speech recognition by automatically extracting features and performing end-to-end training. Today, deep learning has become the core technology of speech recognition and is widely used in smart assistants, speech-to-text, speech translation and other fields. In terms of speech emotion recognition, the advantages of deep learning are also significant. Traditional emotion recognition methods mostly rely on manually extracted features, such as pitch, volume, rhythm, etc. Although they can provide certain information for emotion classification, they are unable to cope with complex emotions. The emotion recognition system based on deep learning, especially in combination with CNN, LSTM and other network structures, can automatically learn richer emotional features from a large amount of speech data, improving the accuracy and robustness of emotion recognition. This has enabled speech emotion recognition to achieve breakthrough progress in multiple application scenarios, especially in the fields of intelligent transportation, customer service, mental health, etc., showing great application potential.
@@ 2.2 Conformal Prediction
CP Angelopoulos and Bates (2021); Ye et al. (2024); Barber et al. (2023); Wang et al. (2024); Sun and Yu (2023); Wang et al. (2025b) is an uncertainty quantification method that provides strict statistical guarantees for machine learning models. Unlike traditional point prediction, CP provides a set of possible prediction labels for each input sample, rather than just a single prediction label, by generating a prediction set containing the prediction results. The core idea is to calculate the nonconformity score to measure the difference between each predicted label and the model’s prediction result, thereby deciding whether the label should be included in the prediction set. The nonconformity score measures the model’s confidence in the prediction result. For classification tasks, it is based on the difference between the model’s predicted probability of the true category and the actual category; for regression tasks, it is the difference between the predicted value and the true value. Before generating the prediction set, Conformal Prediction first uses a calibration set to estimate the distribution of the nonconformity score. The calibration set is a part of the data separated from the training set and is specifically used to measure the prediction uncertainty of the model. By calculating the nonconformity score in the calibration set, Conformal Prediction generates a quantitative value that defines which prediction labels the model should include at a given confidence level. In the test phase, for each new sample, the model outputs a predicted label and its corresponding nonconformity score. Then, CP generates a prediction set by comparing the inconsistency score of each label with the quantized value in the calibration set. The prediction set contains all labels whose inconsistency scores are less than or equal to the quantized value. In this way, the generated prediction set can ensure that the true label always appears in the prediction set at the specified confidence level. This provides a coverage guarantee, that is, at the set confidence level, the probability that the generated prediction set contains the true label is at least the confidence level (such as 90%). This guarantee applies to all models and does not depend on any assumed data distribution, so it is called a distribution-independent prediction method.
@@ 3.1.2 Mathematical Proof Supplement
3.2 Risk Control in speech emotion recognition task1
To adapt the framework of split conformal prediction to the Speech emotion recognition task for user - specified guarantees of task - specific performance, we define a loss function for each calibration data, formulated as
@@ 3.2 Risk Control in speech emotion recognition task1
Figure 1: Conformal Risk Control Framework
Conformal Risk Control (CRC) provides statistical guarantees for high - risk classification tasks through the exchangeable data assumption and custom - defined loss functions. Given $N$ calibration data points and 1 test data point (with a total of $K$ classes), CRC first constructs a prediction set for the sample $x_{i}$:
@@
3.3 Comparison of two frameworks
Conformal Risk Control has a core drawback compared to Split Conformal Prediction in terms of lower computational efficiency, mainly due to its need for multiple traversals of the calibration data to optimize the threshold. Specifically:
The threshold calculation of SCP only requires a single sorting of the non-conformity scores $s_{t}=1-\hat{f}(X_{t})Y^{*}_{t}$ on the calibration set, and taking the quantile $\hat{q}$, with a time complexity of $O(N\log N)$
CRC needs to traverse candidate thresholds $\beta$. For each candidate value, it calculates the average loss on the calibration set $L_{N}(\beta)=\frac{1}{N}\sum_{i=1}^{N}l_{i}$, and then verifies the condition $\frac{NL_{N}(\beta)+B}{N+1}\leq\alpha$. Assuming the number of candidate $\beta$ is $M$, its time complexity is $O(M\cdot N)$ Even with binary search optimization (iteration times $O(\log M)$), it still requires multiple complete traversals of the calibration set, leading to a significant increase in computational load.
This efficiency bottleneck is particularly prominent in large-scale calibration sets ($N\gg 1$) or high-real-time scenarios (such as autonomous driving), while the lightweight single sorting of SCP has more advantages. The flexibility of CRC comes at the cost of efficiency, and one needs to balance the choice according to task requirements.
3.4 Mini-batch conformal prediction theory under non-IID1
In light of the limitations of traditional conformal prediction, which relies on global data exchangeability Barber et al. (2023); Kang et al. (2024); Wang et al. (2025c), we propose the assumption of within - batch local exchangeability. We prove that even when there are differences in data distributions between batches (i.e., batch - to - batch exchangeability may not hold), we can still strictly guarantee the global coverage probability by constructing a non - negative test supermartingale with a forcing operation. This framework breaks through the strong dependence of classical theories on independent and identically distributed data, providing universal theoretical support for small - batch scenarios such as time - series data and online learning.
E - value Gauthier et al. (2025), as a non - negative random variable satisfying $E[E]\leq 1$, can directly construct conformal sets through probability inequalities, offering a more flexible uncertainty quantification tool. For the calibration set $D_{cal}=\{(X_{i},Y_{i})\}_{i=1}^{n}$ and the test sample $(X_{n+1},Y_{n+1})$, the E - value is defined as:
$E=\frac{S_{n+1}}{\frac{1}{n+1}\sum_{i=1}^{n+1}S_{i}}$ (12)
$M_{t}=\max\{(1-\gamma)E_{t}^{m_{-}true}+\gamma M_{t-1},1\}$ (14)
where
$E_{t}^{m_{-}true}=\frac{(n_{t}+1)\times s_{n_{t}+1}^{m_{-}true}}{\sum_{i=1}^{n_{t}}s_{i}+s_{n_{t}+1}^{m_{-}true}}$ (15)
  • Measurability: To determine whether $M_{t}$ is measurable after defining the filtration, we use mathematical induction. When $t=0$, $M_{0}=1$. Since a constant function is measurable with respect to any $\sigma$ algebra, $M_{0}$ is $F_{0}$ measurable, where $F_{0}=\{\varnothing,\Omega\}$ is the trivial $\sigma$ algebra. Assume that $M_{t-1}$ is $F_{t-1}$ measurable. Because $F_{t-1}\subseteq F_{t}$, $M_{t-1}$ is also $F_{t}$ measurable. For the $t$ batch, $E_{t}^{m_{-}true}$ only depends on the non - conformity scores $S_{1}^{t},\cdots,S_{n_{t}}^{t},S_{n_{t}+1}^{t}$ of the $t$ th batch, and these scores are all in $F_{t}$, so $E_{t}^{m_{-}true}$ is $F_{t}$ measurable. According to the update rule, let $Z_{t}=(1-\gamma)E_{t}^{m_{-}true}+\gamma M_{t-1}$. Since $E_{t}^{m_{-}true}$ and $M_{t-1}$ are both $F_{t}$ measurable, the linear combination $Z_{t}$ is also $F_{t}$ measurable. Let $A_{t}=\{\omega\in\Omega:Z_{t}(\omega)<1\}$. Since $Z_{t}$ is $F_{t}$ measurable, $A_{t}\in F_{t}$. We can rewrite $M_{t}$ as $M_{t}=1\cdot I_{A_{t}}+Z_{t}\cdot I_{A_{t}^{c}}$, where $I_{A_{t}}$ is the indicator function of the set $A_{t}$. The indicator functions $I_{A_{t}}$ and $I_{A_{t}^{c}}$ are both $F_{t}$ measurable, and the products of the constant 1 and the measurable function $Z_{t}$ with the indicator functions are also $F_{t}$ measurable. Thus, their sum $M_{t}$ is $F_{t}$ measurable. In conclusion, by mathematical induction, we have proved that for the defined filtration $\{F_{t}\}_{t\geq 0}$, $M_{t}$ is $F_{t}$ measurable.
  • Decreasing conditional expectation: Under the assumption of within - batch data exchangeability, $E[E_{t}^{m_{-}true}|F_{t-1}]=1$. Then the conditional expectation of the supermartingale is
    $E[M_{t}|F_{t-1}]=(1-\lambda)E[E_{t}^{m_{-}true}|F_{t-1}]+\lambda M_{t-1}=(1-\lambda)+\lambda M_{t-1}$ (16)
$\hat{C}(x_{t})=\left\{c\in Y:\max\{(1-\gamma)\frac{(n_{t}+1)\cdot s(x_{t},c)}{\sum_{i=1}^{n_{t}}s(x_{t},y_{i})+s(x_{t},c)}+\gamma M_{t-1},1\}<\frac{1}{\alpha}\right\}$ (17)
Algorithm 1 Adaptive Non-Conformal Prediction with Supermartingale (ANCP-S)

### v4

@@ Abstract
Road rage, often triggered by emotional suppression and sudden outbursts, significantly threatens road safety by causing collisions and aggressive behavior. Speech emotion recognition technologies can mitigate this risk by identifying negative emotions early and issuing timely alerts. However, current SER methods, such as those based on hidden markov models and Long short-term memory networks, primarily handle one-dimensional signals, frequently experience overfitting, and lack calibration, limiting their safety-critical effectiveness. We propose a novel risk-controlled prediction framework providing statistically rigorous guarantees on prediction accuracy. This approach employs a calibration set to define a binary loss function indicating whether the true label is included in the prediction set. Using a data-driven threshold $\beta$, we optimize a joint loss function to maintain an expected test loss bounded by a user-specified risk level $\alpha$. Evaluations across six baseline models and two benchmark datasets demonstrate our framework consistently achieves a minimum coverage of $1-\alpha$, effectively controlling marginal error rates despite varying calibration-test split ratios (e.g., 0.1). The robustness and generalizability of the framework are further validated through an extension to small-batch online calibration under a local exchangeability assumption. We construct a non-negative test martingale to maintain prediction validity even in dynamic and non-exchangeable environments. Cross-dataset tests confirm our method’s ability to uphold reliable statistical guarantees in realistic, evolving data scenarios.
@@ 1 Introduction
In recent years, road rage has emerged as a growing global traffic safety concern. With rapid urbanization and increasing vehicle ownership, heightened traffic congestion and driving stress have led to a rise in extreme emotional responses among drivers. Empirical studies indicate that road rage not only contributes significantly to traffic accidents but also intensifies conflicts between drivers and other road users, thereby posing a serious threat to public safety. According to the National Highway Traffic Safety Administration (NHTSA) [1, 2], over one-third of traffic accidents are associated with emotional loss of control by drivers—a proportion that continues to rise.
Speech Emotion Recognition technology [3, 4], which identifies emotional states through the analysis of vocal signals, has increasingly been recognized as an effective approach to mitigating road rage and other emotion-induced driving risks. [5] By capturing vocal characteristics in real time and inferring the driver’s emotional state, SER enables timely feedback to intelligent traffic management systems, facilitating the early detection and intervention of potentially hazardous driving behaviors. Beyond its applications in smart vehicles for precise emotion monitoring, SER has also shown promise in broader emotion-centric domains such as customer service and mental health assessment [6, 7]. Given its cross-disciplinary relevance, SER is rapidly gaining traction as a core research focus in intelligent transportation systems, autonomous driving technologies, and affective computing.
With the rapid advancement of deep learning, speech emotion recognition based on convolutional neural networks (CNNs) has achieved notable success. In this study, we utilize Mel-spectrogram features to represent speech signals and differentiate among emotional categories through comparative experiments across models. Despite these advancements, CNN-based SER models remain susceptible to overfitting during training, leading to degraded generalization on previously unseen data. [8, 9] Moreover, most deep learning models lack inherent mechanisms to quantify predictive uncertainty, making their outputs potentially unreliable in real-world applications.
Figure 1: Overall framework.
Recent research has highlighted the issue of miscalibration: Although CNNs often achieve high classification accuracy in SER tasks, their prediction confidence tends to be overestimated. For instance, Zhao et al. [10] demonstrated that this overconfidence substantially increases the risk of misclassification, particularly for low-confidence samples. Thus, improving the uncertainty quantification of SER models has become essential to enhancing their reliability and robustness [11].
To address these challenges, Conformal Prediction (CP) provides a statistically rigorous framework for uncertainty quantification and prediction calibration [12]. CP constructs prediction sets that are guaranteed to contain the true label with a user-defined confidence level, thus enhancing model trustworthiness in safety-critical applications. Our overall framework is illustrated in Figure 1. In the context of SER [13], we extract Mel-spectrogram features and train five neural network architectures with comparable parameter sizes on the IEMOCAP dataset as base models. Following model training, the CP procedure is applied in three steps: (1) defining nonconformity scores for each prediction; (2) computing a calibration quantile $\hat{q}$ from a held-out calibration set based on a user-specified risk level $\alpha$; and (3) generating prediction sets for test samples that satisfy the marginal coverage guarantee. By outputting emotion label sets with formal confidence guarantees, our approach establishes a principled statistical foundation for improving the reliability of speech emotion recognition systems.
While conformal prediction provides formal coverage guarantees for prediction sets, it primarily focuses on marginal coverage and does not account for task-specific performance criteria. To address this limitation, the risk-controlled conformal prediction (RCCP) framework [14] extends traditional CP by bounding the expected value of arbitrary monotonic loss functions, thereby enabling principled risk management in complex real-world tasks. This generalization maintains statistical rigor while offering flexibility in defining application-specific loss metrics, making the approach broadly applicable across diverse domains. By combining CP with risk control, the framework preserves coverage guarantees while aligning predictions with user-defined performance requirements. However, conventional CP methods assume exchangeability among data points—a condition that holds in controlled environments but is frequently violated in real-world scenarios, such as speech emotion recognition, where data distributions often shift across speakers, sessions, or acoustic conditions. In such non-exchangeable settings, standard CP methods may fail to uphold their theoretical guarantees, thus compromising reliability. To address this challenge, we propose a novel framework that extends CP to dynamic, non-exchangeable environments via a mini-batch online calibration mechanism. Specifically, we relax the global exchangeability assumption by introducing a local exchangeability assumption within individual mini-batches. By constructing a non-negative test supermartingale, our method adaptively calibrates prediction sets in an online manner, ensuring valid coverage even under distributional shifts between batches—such as variations in speaker identity or environmental noise. This framework provides a statistically sound and practically robust solution for deploying CP-based uncertainty quantification in dynamic real-world SER applications.
To validate the effectiveness of our framework, we conducted a series of empirical studies. Preliminary evaluations on the TESS dataset compared multiple feature extraction methods, with Mel-spectrograms demonstrating superior performance and thus selected for downstream tasks. We then trained several neural network architectures (e.g., ResNet, MobileNetV3, ShuffleNetV2) on IEMOCAP, where a clear overfitting problem emerged, despite perfect training accuracy, test performance degraded significantly. To ensure prediction reliability, we integrated the CP framework, conducting both within-dataset and cross-dataset experiments. Results confirmed that CP consistently maintained desired coverage levels, even under data distribution shifts, highlighting its robustness. Moreover, we observed that prediction set size serves as an effective indicator of model uncertainty [15], prompting the use of a risk-controlled loss function to adaptively regulate prediction confidence. Finally, to address non-exchangeable settings, we introduced a mini-batch martingale-based calibration strategy. Experiments with mismatched calibration and test datasets (IEMOCAP to TESS) demonstrated that, unlike standard CP, our approach preserved high coverage under local exchangeability assumptions, confirming its practical utility in dynamic environments.
@@ 2.1 Speech Recognition
The evolution of speech recognition technology has progressed from early template matching systems in the 1950s to modern deep learning approaches. Traditional models, such as hidden Markov models and Gaussian mixture models (GMMs), enabled statistical modeling of speech sequences but struggled with noise and variability [16, 17]. The advent of deep learning transformed the field: convolutional and recurrent neural networks, including long short-term memory architectures, enabled end-to-end training and improved robustness in real-world scenarios [18, 19]. In the context of speech emotion recognition, deep learning has similarly replaced hand-crafted feature engineering with automatic representation learning. [20, 21] As reviewed by Schuller [22], hybrid CNN-LSTM models now form the backbone of state-of-the-art SER systems, capable of capturing complex emotional nuances in speech data. Today, deep learning stands as the foundation of both speech recognition and emotion understanding, powering a wide range of applications from intelligent assistants to mental health monitoring.
@@ 2.2 Conformal Prediction
Conformal Prediction  [23, 24, 25, 26, 27, 28] is an uncertainty quantification framework that offers finite-sample statistical guarantees for predictive models. Unlike traditional point predictions, CP outputs a prediction set—a subset of possible labels—that is guaranteed to contain the true label with a user-specified confidence level (e.g., 90%). The core idea of CP is to compute a nonconformity score that quantifies how unusual a predicted label is, given the model’s behavior. For classification tasks, this score is often based on the model’s confidence in the correct class (e.g., 1 minus the predicted probability of the true label); for regression, it typically corresponds to the absolute error between predicted and true values. During calibration, CP uses a held-out calibration set to estimate the distribution of nonconformity scores. A quantile threshold is then determined based on the desired coverage level. At test time, for each new input, the model evaluates nonconformity scores across candidate labels, and the prediction set includes all labels with scores below the calibrated threshold. Importantly, CP makes no assumptions about the underlying data distribution, requiring only that the calibration and test samples are exchangeable. This distribution-free property enables CP to provide reliable coverage guarantees across a wide range of models and applications.
@@ 3 Method
Figure 2: Illustration of the Risk-controlled Conformal Prediction Framework
@@ 3.1.2 Mathematical Proof Supplement
3.2 Risk-controlled Conformal Prediction in speech emotion recognition task
To adapt the framework of split conformal prediction to the speech emotion recognition task for user - specified guarantees of task - specific performance,we have developed a risk-controlled conformal prediction framework, as shown in the figure 2, which explains how the framework works, we define a loss function for each calibration data, formulated as
@@ 3.2 Risk-controlled Conformal Prediction in speech emotion recognition task
RCCP provides statistical guarantees for high - risk classification tasks through the exchangeable data assumption and custom - defined loss functions. Given $N$ calibration data points and 1 test data point (with a total of $K$ classes), RCCP first constructs a prediction set for the sample $x_{i}$:
@@
Algorithm 1 Conformal Prediction with Nonnegative Test Supermartingale
@@
3.3 Comparison of two frameworks
RCCP has a core drawback compared to SCP in terms of lower computational efficiency, mainly due to its need for multiple traversals of the calibration data to optimize the threshold. Specifically: The threshold calculation of SCP only requires a single sorting of the non-conformity scores $s_{t}=1-\hat{f}(X_{t})_{Y^{*}_{t}}$ on the calibration set, and taking the quantile $\hat{q}$, with a time complexity of $O(N\log N)$. RCCP needs to traverse candidate thresholds $\beta$. For each candidate value, it calculates the average loss on the calibration set $L_{N}(\beta)=\frac{1}{N}\sum_{i=1}^{N}l_{i}$, and then verifies the condition $\frac{NL_{N}(\beta)+B}{N+1}\leq\alpha$. Assuming the number of candidate $\beta$ is $M$, its time complexity is $O(M\cdot N)$ Even with binary search optimization (iteration times $O(\log M)$), it still requires multiple complete traversals of the calibration set, leading to a significant increase in computational load. This efficiency bottleneck is particularly prominent in large-scale calibration sets ($N\gg 1$) or high-real-time scenarios (such as autonomous driving), while the lightweight single sorting of SCP has more advantages. The flexibility of RCCP comes at the cost of efficiency, and one needs to balance the choice according to task requirements.
3.4 Mini-batch conformal prediction theory under Non-exchangeable Data
In light of the limitations of traditional conformal prediction, which relies on global data exchangeability [25, 29, 30], we propose the assumption of within - batch local exchangeability. We prove that even when there are differences in data distributions between batches (i.e., batch - to - batch exchangeability may not hold), we can still strictly guarantee the global coverage probability by constructing a non - negative test supermartingale with a forcing operation. The framework is shown in the algorithm 1. This framework breaks through the strong dependence of classical theories on global data exchangeability, providing universal theoretical support for small - batch scenarios such as time - series data and online learning.
E - value [31, 32], as a non - negative random variable satisfying $E[E]\leq 1$, can directly construct conformal sets through probability inequalities, offering a more flexible uncertainty quantification tool. For the calibration set $D_{cal}=\{(X_{i},Y_{i})\}_{i=1}^{n}$ and the test sample $(X_{n+1},Y_{n+1})$, the E - value is defined as:
$E=\frac{s_{n+1}}{\frac{1}{n+1}\sum_{i=1}^{n+1}s_{i}}$ (12)
$M_{t}=\max\{(1-\gamma)E_{t}+\gamma M_{t-1},1\}$ (14)
where
$E_{t}=\frac{(n_{t}+1)\times s_{n_{t}+1}}{\sum_{i=1}^{n_{t}}s_{i}+s_{n_{t}+1}}$ (15)
  • Measurability: To determine whether $M_{t}$ is measurable after defining the filtration, we use mathematical induction. When $t=0$, $M_{0}=1$. Since a constant function is measurable with respect to any $\sigma$ algebra, $M_{0}$ is $F_{0}$ measurable, where $F_{0}=\{\varnothing,\Omega\}$ is the trivial $\sigma$ algebra. Assume that $M_{t-1}$ is $F_{t-1}$ measurable. Because $F_{t-1}\subseteq F_{t}$, $M_{t-1}$ is also $F_{t}$ measurable. For the $t$ batch, $E_{t}$ only depends on the non - conformity scores $s_{1}^{t},\cdots,s_{n_{t}}^{t},s_{n_{t}+1}^{t}$ of the $t$ th batch, and these scores are all in $F_{t}$, so $E_{t}^{m_{-}true}$ is $F_{t}$ measurable. According to the update rule, let $Z_{t}=(1-\gamma)E_{t}+\gamma M_{t-1}$. Since $E_{t}$ and $M_{t-1}$ are both $F_{t}$ measurable, the linear combination $Z_{t}$ is also $F_{t}$ measurable. Let $A_{t}=\{\omega\in\Omega:Z_{t}(\omega)<1\}$. Since $Z_{t}$ is $F_{t}$ measurable, $A_{t}\in F_{t}$. We can rewrite $M_{t}$ as $M_{t}=1\cdot I_{A_{t}}+Z_{t}\cdot I_{A_{t}^{c}}$, where $I_{A_{t}}$ is the indicator function of the set $A_{t}$. The indicator functions $I_{A_{t}}$ and $I_{A_{t}^{c}}$ are both $F_{t}$ measurable, and the products of the constant 1 and the measurable function $Z_{t}$ with the indicator functions are also $F_{t}$ measurable. Thus, their sum $M_{t}$ is $F_{t}$ measurable. In conclusion, by mathematical induction, we have proved that for the defined filtration $\{F_{t}\}_{t\geq 0}$, $M_{t}$ is $F_{t}$ measurable.
  • Decreasing conditional expectation: Under the assumption of within - batch data exchangeability, $E[E_{t}|F_{t-1}]=1$. Then the conditional expectation of the supermartingale is
    $E[M_{t}|F_{t-1}]=(1-\lambda)E[E_{t}|F_{t-1}]+\lambda M_{t-1}=(1-\lambda)+\lambda M_{t-1}$ (16)
$\hat{C}(x_{t})=\left\{m\in Y:\max\{(1-\gamma)\frac{(n_{t}+1)\cdot s(x_{t},m)}{\sum_{i=1}^{n_{t}}s_{i}+s(x_{t},m)}+\gamma M_{t-1},1\}<\frac{1}{\alpha}\right\}$ (17)
4 Experiments
4.1 Experimental Settings